# Proactive Intra-GPU Disaggregation of Prefill and Decode in LLM Serving

### v4

@@ 3.3. Limitation of Static Partitioning
Section 3.2 shows that prefill and decode exhibit diminishing returns with increasing SM allocation, motivating phase-aware compute partitioning. However, even a well-chosen static partition is insufficient in practice. This is because compute demand, while often estimable from inputs like chunk size or sequence length, does not account for dynamic memory behavior during serving. Prefill and decode phases contend for memory bandwidth in ways that depend on runtime factors such as evolving KV-cache sizes, prompt distributions, and decode lengths.
To illustrate this, we co-execute prefill chunks [^2] Following Sarathi-Serve (Agrawal et al., 2024), long prompts are split into smaller segments and processed alongside decode requests. These prefill chunks actively read and write to the KV cache. with pure decode batches under a fixed SM partition. As shown in Figure 6(a), increasing the prefill KV length from 2,000 to 10,000 increases decode latency by 36%, despite the decode workload itself remaining constant. This slowdown stems from bandwidth contention: prefill attention layers issue large KV reads that overlap with latency-critical memory accesses from the decode phase.
Figure 6(b) further shows that prefill memory traffic is highly irregular and fluctuates significantly over time, making contention both unpredictable and workload-dependent.
These results highlight a fundamental limitation: even when compute needs are analyzable, memory bandwidth pressure can vary rapidly due to asynchronous phase behavior. Static SM partitioning cannot react to this dynamic interference, leading to resource underutilization and inflated latency.
{mdframed}
[backgroundcolor=gray!7, linewidth=0.8pt] Insight 2. Intra-GPU PD disaggregation with static resource allocation is insufficient under dynamic memory contention. Effective intra-GPU scheduling requires runtime-aware, fine-grained SM reallocation that adapts to emergent phase interference.
@@ 4. Design
We present Nexus, a lightweight monolithic LLM serving system that enables intra-GPU prefill–decode disaggregation through proactive SM partitioning and phase-aware scheduling. Unlike reactive systems that adjust only after SLO violations, Nexus proactively predicts performance bottlenecks using a contention-aware cost model and dynamically reallocates resources in real time. This enables high utilization without introducing interference. As shown in Figure 7, Nexus introduces three core mechanisms to enable this fine-grained separation:
@@ 4.1. Dynamic SM Partitioning via Cost Model
Nexus’s cost model estimates the latency of prefill and decode under SM allocations without execution. This avoids reliance on feedback-based tuning or runtime latency profiling, and supports proactive, per-batch optimization not possible in prior reactive systems.
@@
Compute Latency. We estimate the compute latency of each operator $o\in\texttt{PrefillOps}\cup\texttt{DecodeOps}$ based on its FLOP count $c_{o}$ and the SM ratio $r$ assigned to its stage. While latency ideally scales inversely with compute share ($1/r$), real-world performance can deviate depending on the kernel(§3.2).
@@
These coefficients ($R_{\text{sat}},\lambda$) are extracted from a one-time profiling pass per model, avoiding the need for runtime latency fitting and transferable across workloads.
Memory Contention Aware Model. As shown in §3.3, memory contention can significantly affect latency when prefill and decode execute concurrently on shared hardware. Unlike prior work that ignores phase overlap or assumes fixed latency models, we model decode attention latency as a function of runtime overlap with memory-bound prefill activity and the relative memory bandwidth during contention.
@@
Assuming full bandwidth saturation during each overlap window, we allocate effective bandwidth to decode based on its share of memory traffic. Let $m_{d}$ be the total memory bytes accessed by decode attention, $m_{p1}$ the bytes accessed by prefill attention, and $m_{p2}$ those accessed by prefill’s dense operators. [^3] We estimate memory access volume for linear layers using known parameter sizes from the model architecture. For attention, we weigth memory access and calculate KV memory traffic based on the number of cached tokens tracked by vLLM, multiplied by the per-token key/value size determined from model dimensions. Then the effective bandwidth for decode is:
@@
Given the cost model, Nexus selects an SM partition that balances performance and memory pressure. However, since prefill and decode run concurrently and compete for GPU compute, GPU memory and GPU memory bandwidth, optimizing both simultaneously can lead to inefficient tradeoffs.
To resolve this, Nexus formulates a dual-objective latency optimization: prioritize one phase while constraining the other to remain within a slowdown budget. This allows flexible tradeoffs between TBT and TTFT, depending on workload state. The choice is also guided by runtime signals, such as KV cache usage, to avoid pathological behaviors like OOM or long waiting time.
@@
Here, $T^{\text{min}}$ denotes the ideal latency when a stage is allocated all SMs, and $\alpha,\beta>1$ are slack variables controlling tolerable slowdowns in the non-prioritized stage. Unlike SLO thresholds, these slack constraints are batch-dependent and derived from the cost model itself.
Runtime Switching. We select the objective mode based on live KV cache usage $KV_{u}$: when $KV_{u}$ is low, prioritizing prefill improves TTFT by accelerating prompt ingestion. When $KV_{u}$ is high, prioritizing decode helps reduce memory pressure by completing generation and evicting old KV entries. This feedback mechanism enables resource allocation to respond to workload phases and memory constraints.
Objective $\text{Objective Mode}=\begin{cases}\text{Prefill-prioritized}&\text{if }KV_{u}\leq KV_{\text{switch}}\\ \text{Decode-prioritized}&\text{otherwise}\end{cases}$
@@
9$\triangleright$ Buffer zone check to suppress unstable or small changes
@@ 22    $R\leftarrow R_{\text{cur}}$
23$\triangleright$ Phase 1: Decrease until constraint is satisfied
@@ 25         $R\leftarrow R-1$
27$\triangleright$ Phase 2: Increase target share until constraint is at limit
@@ 33         $R\leftarrow R+1$
35    return (target is prefill? $(R,100-R)$ : $(100-R,R)$)
@@ 4.1.3. Greedy SM Search
Efficiency. We avoid global constrained solvers due to their overhead. Instead, a simple greedy search provides robustness and transparency while converging in just 2–4 iterations. This efficiency is also critical for per-batch decision making during inference.
@@ 4.2. Hysteresis-Based SM Repartitioning for Stability
This simple mechanism retains the adaptability of fine-grained decisions while avoiding the instability of overreactive switching. Alternative approaches like naively reducing the update frequency sacrifices responsiveness, while large step sizes risk overshooting optimal ratios. In contrast, our hysteresis design filters out noise and enables smooth SM reallocation under dynamic workloads.
@@
[…]
6         $r.\text{score}\leftarrow r.\text{remaining}-\gamma\cdot r.\text{age}$triangleright Antistarvation
@@ 4.3. Phase-Specific Scheduler
Since Nexus separates prefill and decode into concurrent batches, we can further exploit this with phase-specific optimizations. While SM partitioning governs how much resource each phase gets, our schedulers govern how that resource is consumed. Particularily, we employ scheduler policies tailored to the latency and compute characteristics of each phase.
@@ 4.3.2. Decode Scheduler
The decode phase controls TBT. Unlike prefill, decode scheduling operates at a finer granularity, with each request contributing only a single token to the active batch. Thus, we adopt a simple First-Come-First-Serve (FCFS) policy. FCFS ensures fairness, incurs minimal overhead, and avoids token-level starvation. While more sophisticated schemes that consider context window size or memory bandwidth are theoretically possible, we find that their impact is limited.
@@ 6.1. Experimental Setup
Baselines. We compare Nexus against four representative LLM serving engines, all configured with the same tensor parallelism, chunked prefill, and scheduling budget for fair comparison. We plan to integrate Semi-PD in a future update to provide a more complete intra-GPU PD disaggregation system comparison.
@@ 6.2.1. End-to-End Single-GPU Performance
TTFT. As shown in Figure 9 (columns 3–4), Nexus achieves the lowest or near-lowest average TTFT across all workloads. It improves TTFT by 2–20$\times$ over vLLM and up to 1.6$\times$ over SGLang through SPF scheduling and dynamic SM reallocating. The latter reallocates GPU resources at runtime to reduce prefill–decode contention, further amplifying SPF’s benefits. vLLM and SGLang suffer from head-of-line blocking under FCFS, though SGLang fares better due to its optimized runtime. FastServe reduces average TTFT via skip-join MLFQ, but hurts P95 due to deprioritizing long prompts. Compared to vLLM-P/D, which avoids contention by separating phases across GPUs, Nexus matches its TTFT in Mixed Workload and remains within 10% on Long Data Collections and Arxiv while using a single GPU.
To ensure fairness, Nexus includes a tunable anti-starvation mechanism within its prefill scheduler. Under current settings, it improves P95 TTFT by 2–3$\times$ over vanilla vLLM and narrows the gap with SGLang and vLLM-P/D in Long Data Collections and Arxiv Workloads, while maintaining consistent advantages in average latency. In Mixed Workload, Nexus shows worse tail TTFT due to high prompt length diversity, which increases batching variability and makes it harder to protect long requests without hurting throughput.
@@
FastServe degrades sharply under load as it needs to fall back on recomputation. vLLM, which co-schedules prefill and decode, suffers from intra-batch interference, trailing Nexus by 1.24$\times$–1.48$\times$. SGLang improves over vLLM via Radix Attention and runtime optimizations, closely matching Nexus in Long Data Collections Workload, slightly surpassing in Arxiv Workload, but falling behind in Mixed Workload where prompt diversity intensifies decode imbalance.
@@
From Figure 9 (columns 1–2), Nexus consistently delivers the highest throughput among single-GPU systems. In Long Data Collections and Arxiv Workloads, it achieves 1.5–1.8$\times$ higher throughput than vLLM and 1.18-1.27$\times$ higher than SGLang, reflecting more efficient resource scheduling under uniform or moderately variable request patterns.
In Mixed Workload, where prompt diversity and scheduling imbalance are most pronounced, Nexus demonstrates its largest gain: it achieves 1.9$\times$ higher throughput than vLLM, 1.8$\times$ over SGLang, and even 1.4$\times$ over vLLM-P/D, despite the latter using two GPUs and full-phase disaggregation. This is driven by Nexus’s head-light design which aggressively prioritizing short requests and adapting GPU resource allocation dynamically, enabling it to avoid contention and return early outputs with minimal delay.
@@ 6.2.2. End-to-End Multi-GPU Performance
As shown in FIgure 10, Nexus achieves the highest throughput, 2.2$\times$ over vLLM and 2$\times$ over SGLang, while using the same hardware. These gains come from efficient intra-engine phase separation and the ability to maintain high concurrency without overwhelming shared compute or memory resources.
More importantly, this throughput does not come at the cost of latency. Nexus delivers 2–3$\times$ lower average TTFT and 1.5–2$\times$ lower TBT than vLLM and SGLang, showing strong responsiveness across both prefill and decode phases. At the tail, P95 TTFT is slightly higher than SGLang but matches vLLM, while P95 TBT is nearly identical across all systems.
@@ 6.4. Latency Breakdown
Queuing Delay. Waiting time dominates total latency under load, and here Nexus demonstrates its greatest advantage. In Long Data Collections, Nexus reduces waiting time by 4$\times$ over vLLM and 2$\times$ over SGLang. In Mixed Workload, which involves greater request variability, it improves further, achieving 5$\times$ lower wait time than monolithic baselines, and 2$\times$ less than vLLM-P/D. These gains stem from Nexus’s effective shortest-prompt-first scheduling and adaptive GPU resource allocation, allowing it to maintain concurrency without overloading shared resources.
@@ 7. Related Work
Monolithic LLM Serving Systems. Early LLM serving engines focus on maximizing throughput and memory efficiency within a unified execution engine. Orca (Yu et al., 2022) introduces continuous batching to reduce head-of-line blocking. vLLM (Kwon et al., 2023) eliminates KV cache fragmentation via PagedAttention, while SGLang (Zheng et al., 2023) reduces memory usage through RadixAttention. SarathiServe (Agrawal et al., 2024) mixes chunks of prefill and decode requests to better utilize GPU resources. However, all of these designs treat prefill and decode as indistinguishable units within a shared queue. In contrast, Nexus decouples the two phases at the batching level, enabling independent, phase-specific execution and scheduling while preserving compatibility with existing attention mechanisms.
@@
GPU Multiplexing. Work on GPU multiplexing explores spatial or temporal sharing across models or processes. Systems like GPUlet, Orion, REEF, and Bless (Choi et al., 2022; Strati et al., 2024; Han et al., 2022; Zhang et al., 2025) optimize for small models via fine-grained temporal sharing. MuxServe (Duan et al., 2024) proposes spatial-temporal multiplexing to efficiently serve multiple LLMs, while NanoFlow (Zhu et al., 2024) and Liger (Du et al., 2024) introduce kernel level parallelism that enables concurrent execution of compute bound, network bound, and memory bound operations. Unlike these systems, which focus on generic sharing or inter-model concurrency, Nexus targets intra-GPU, phase-level disaggregation. It dynamically and proactively reallocates GPU resources between prefill and decode phases within a single LLM engine, enabling fine-grained, workload-aware partitioning guided by lightweight performance modeling and coordinated scheduling.
LLM Scheduling Frameworks. Prior works also propose sophisticated schedulers to balance latency and throughput. FastServe (Wu et al., 2023) uses MLFQ with skip-join to avoid prompt variance stalls. VTC (Sheng et al., 2024) and QLM (Patke et al., 2024) target fairness and SLO adherence. Llumnix (Sun et al., 2024) accelerates dynamic scaling via migration-aware scheduling. LightLLM (Gong et al., 2025) and Preble (Srivatsa et al., 2024) improve memory reuse via future prediction or prompt sharing. Yet, all treat each request as a monolithic scheduling unit. Nexus introduces dual schedulers for prefill and decode, each tailored to its phase’s latency sensitivity and compute intensity. This decoupling enables tighter queue control and more efficient GPU resource utilization.
@@ 8. Conclusion
While the asymmetric demands of prefill and decode in LLM inference are well known, their combined impact on intra-GPU contention under dynamic workloads remains underexplored. We show that static isolation and reactive control fall short in managing compute saturation, memory and memory bandwidth contention. Nexus addresses this via a lightweight analytical model, dual-objective optimization, and phase-specific schedulers for proactive, sub-second resource reallocation. Across diverse models and workloads, Nexus improves both throughput and latency, outperforming monolithic and disaggregated baselines with fewer resources.
9. Acknowledgement
We sincerely thank Zikun Li, Xinhao Cheng, Mingyu Li, Linsong Guo, Lifan Wu, Xupeng Miao, Shulai Zhang, Zihao Ye, Xiyuan Cui, Zhanda Zhu, Zhihao Jia, Xingda Wei and Yubin Xia for their kind help on this paper.

### v5

@@ 3.3. Limitation of Static Partitioning
Section 3.2 shows that prefill and decode phases exhibit sharply diminishing returns with increasing SM allocation, motivating finer-grained partitioning. However, even a well-chosen static SM partition is suboptimal at runtime. This is because compute demand, while often estimable from inputs like chunk size or sequence length, does not capture dynamic memory behavior. Prefill and decode contend for bandwidth in ways that depend on evolving KV cache sizes, prompt distributions, and decode lengths that emerge only at runtime.
To illustrate this, we co-execute chunks of prefill [^2] Following Sarathi-Serve (Agrawal et al., 2024), long prompts are split into smaller segments and executed alongside decode requests. These chunks of prefill tasks read and write to the KV cache. and pure decodes under a fixed SM partition. As we can see from Figure 6(a), increasing the prefill KV length from 2000 to 10000 also increases the latency of the exact same decode batch by 36%. This slowdown stems from memory bandwidth contention: prefill attention layers perform large KV reads, which overlap with the decode stage’s latency-critical memory access.
Moreover, Figure 6(b) shows that prefill memory traffic is highly irregular, fluctuating significantly over time and making contention both difficult to predict and workload dependent.
These highlight a core limitation: even when compute demands are static and analyzable, bandwidth pressure varies at runtime due to asynchronous prefill/decode evolution. Static SM partitioning fails to respond to this variation, leading to avoidable contention and degraded latency.
Insight 3. Logical PD disaggregation with static partitioning is insufficient. To adapt to emergent memory contention and shifting runtime demands, systems must have fine-grained, dynamic SM reallocation.
@@ 4. Design
We present Nexus, a lightweight monolithic LLM serving system that enables intra-engine prefill–decode disaggregation. Unlike prior systems that either co-batch prefill and decode or isolate them across GPUs, Nexus partitions GPU resources dynamically, executing both phases concurrently but independently within a single engine.
As shown in Figure 7, Nexus introduces three core mechanisms to enable this fine-grained separation:
@@ 4.1. Dynamic SM Partitioning via Cost Model
Nexus’s cost model estimates the latency of prefill and decode under SM allocations without execution, enabling rapid exploration of latency tradeoffs.
@@
Compute Latency. We estimate the compute latency of each operator $o\in\texttt{PrefillOps}\cup\texttt{DecodeOps}$ based on its FLOP count $c_{o}$ and the SM ratio $r$ assigned to its stage. While latency ideally scales inversely with compute share ($1/r$), real-world performance can deviates depending on kernel(§3.2).
@@
We extract $R_{\text{sat}}$ and $\lambda$ per operator from end-to-end measurements of the full stage (prefill or decode) under varying SM allocations.
Memory Access Latency. As shown in §3.3, memory contention can significantly affect latency when prefill and decode execute concurrently on shared hardware. To capture this effect, we model decode memory latency as a function of (1) temporal overlap with prefill, and (2) the relative memory bandwidth demands of each phase.
@@
Assuming full bandwidth saturation during each overlap window, we allocate effective bandwidth to decode based on its share of memory traffic. Let $m_{d}$ be the total memory bytes accessed by decode attention, $m_{p1}$ the bytes accessed by prefill attention, and $m_{p2}$ those accessed by prefill’s dense operators. [^3] We estimate memory access volume for linear layers using known parameter sizes from the model architecture. For attention, we calculate KV memory traffic based on the number of cached tokens tracked by vLLM, multiplied by the per-token key/value size determined from model dimensions. Then the effective bandwidth for decode is:
@@
Given the cost model, Nexus selects an SM partition that balances performance and memory pressure. However, since prefill and decode run concurrently and compete for resources, optimizing both simultaneously is infeasible.
To resolve this, Nexus formulates a dual-objective latency optimization: prioritize one phase while constraining the other to remain within a slowdown budget. This allows flexible tradeoffs between TBT and TTFT, depending on workload state. The choice is also guided by runtime signals, such as KV cache usage, to avoid pathological behaviors like OOM.
@@
Here, $T^{\text{min}}$ denotes the ideal latency when a stage is allocated all SMs, and $\alpha,\beta>1$ are slack variables controlling tolerable slowdowns in the non-prioritized stage.
Runtime Switching. We select the objective mode based on live KV cache usage $KV_{u}$: when $KV_{u}$ is low, favoring prefill accelerates prompt ingestion; when $KV_{u}$ is high, prioritizing decode helps reduce memory pressure by completing generations and evicting KV. This feedback mechanism enables resource allocation to respond to workload phases and memory constraints.
$\text{Objective Mode}=\begin{cases}\text{Prefill-prioritized}&\text{if }KV_{u}\leq KV_{\text{switch}}\\ \text{Decode-prioritized}&\text{otherwise}\end{cases}$
@@
[…]
@@ 22  $R\leftarrow R_{\text{cur}}$
[…]
@@ 25    $R\leftarrow R-1$
[…]
@@ 33    $R\leftarrow R+1$
35  return (target is prefill? (R,100R)(R,100-R) : (100R,R)(100-R,R))
@@ 4.1.3. Greedy SM Search
Efficiency. The search typically converges within 2–4 cost model evaluations and imposes negligible latency overhead. This fast feedback loop allows Nexus to adapt to runtime contention and shifting workloads.
@@ 4.2. Hysteresis-Based SM Repartitioning for Stability
This simple mechanism retains the adaptability of fine-grained decisions while avoiding the instability of overreactive switching. Alternative approaches like reducing update frequency or increasing step size proved brittle in practice: the former sacrifices responsiveness; the latter risks overshooting optimal ratios. Buffered switching provides a robust, low-overhead tradeoff.
@@
5    $r.\text{remaining}\leftarrow r.\text{prompt\_len}-r.\text{prefilled\_len}$
6    $r.\text{score}\leftarrow r.\text{remaining}-\gamma\cdot r.\text{age}$ $\triangleright$ Antistarvation
@@ 4.3. Phase-Specific Scheduler
Since Nexus separates prefill and decode into concurrent batches, we can further exploit this with phase-specific optimizations. Particularily, we employee tailored scheduler policies that address the distinct latency and resource profiles.
@@ 4.3.2. Decode Scheduler
The decode phase controls TBT. Unlike prefill, decode scheduling operates at a finer granularity, with each request contributing only a single token to the active batch. Thus, we adopt a simple First-Come-First-Serve (FCFS) policy. FCFS ensures fairness, incurs minimal overhead, and avoids token-level starvation. While more sophisticated schemes that considers context window size or memory bandwidth are theoretically possible, we find that their impact is limited in practice as they are considered in SM partitioning.
@@ 6.1. Experimental Setup
Baselines. We compare Nexus against four representative LLM serving engines, all configured with the same tensor parallelism, chunked prefill, and scheduling budget for fair comparison.
@@ 6.2.1. End-to-End Single-GPU Performance
[…]
@@
[…]
@@
[…]
@@ 6.2.2. End-to-End Multi-GPU Performance
[…]
@@ 6.4. Latency Breakdown
[…]
@@ 7. Related Work
Monolithic LLM Serving Systems. Early LLM serving engines focus on maximizing throughput and memory efficiency within a unified execution model. Orca (Yu et al., 2022) introduces continuous batching to reduce head-of-line blocking. vLLM (Kwon et al., 2023) eliminates KV cache fragmentation via PagedAttention, while SGLang (Zheng et al., 2023) reduces memory usage through RadixAttention. SarathiServe (Agrawal et al., 2024) mixes chunks of prefill and decode requests to better utilize GPU resources. However, all of these designs treat prefill and decode as indistinguishable units within a shared queue. In contrast, Nexus decouples the two phases at the batching level, enabling independent, phase-specific execution and scheduling while preserving compatibility with existing attention mechanisms.
LLM Scheduling Frameworks. Recent works propose sophisticated schedulers to balance latency and throughput. FastServe (Wu et al., 2023) uses MLFQ with skip-join to avoid prompt variance stalls. VTC (Sheng et al., 2024) and QLM (Patke et al., 2024) target fairness and SLO adherence. Llumnix (Sun et al., 2024) accelerates dynamic scaling via migration-aware scheduling. LightLLM (Gong et al., 2025) and Preble (Srivatsa et al., 2024) improve memory reuse via future prediction or prompt sharing. Yet, all treat each request as a monolithic scheduling unit. Nexus introduces dual schedulers for prefill and decode, each tailored to its phase’s latency sensitivity and compute intensity. This decoupling enables tighter queue control and more efficient GPU resource utilization.
@@
Intra-GPU PD Disaggregation Systems. There are some prior works on intra-engine PD disaggregation (Cui et al., 2025; Lin et al., 2025). Drift (Cui et al., 2025) introduces an adaptive gang scheduling mechanism, a contention-free performance model, and an SLO-aware dispatching policy to enable intra-engine prefill–decode separation. Bullet (Lin et al., 2025) proposes a comprehensive system that includes: (1) a performance estimator for building a profile-augmented analytical model; (2) an SLO-aware scheduler to dynamically balance the compute load between prefill and decode phases; and (3) a resource manager capable of delivering fast yet accurate resource configurations. The key differences between their and Nexus are as follows: Nexus employs a contention-based cost model to estimate latency, formulates a dual-objective optimization problem, and uses a greedy search to determine the optimal SM partitioning. In addition, our scheduler is phase-aware and explicitly considers the distinct characteristics of prefill and decode stages, which contrasts with others’ SLO-aware scheduling approach.
@@
GPU Multiplexing. Recent work explores fine-grained GPU sharing via spatial or temporal partitioning. NVIDIA MPS (Nvidia, 2025e) and MIG (Nvidia, 2025d) provide coarse-grained isolation, while GreenContext (Nvidia, 2025b) enables dynamic intra-process SM partitioning. Systems like GPUlet, Orion, REEF, and Bless (Choi et al., 2022; Strati et al., 2024; Han et al., 2022; Zhang et al., 2025) improve utilization for small models through temporal multiplexing. MuxServe (Duan et al., 2024) proposes spatial-temporal multiplexing to efficiently serve multiple LLMs, while NanoFlow (Zhu et al., 2024) and Liger (Du et al., 2024) introduce kernel level parallelism that enables concurrent execution of compute bound, network bound, and memory bound operations. PoD (Kamath et al., 2025) fuses prefill attention and decode attention into one kernel to reduce overhead.
Nexus differs in three key aspects: (1)it disaggregates prefill and decode into different batch and executes these batch concurrently; and (2) it dynamically reallocate SMs across prefill and decode stages in a single engine, adapting to workload shifts; and (3) it introduces intra-engine, phase-specific schedulers for coordinated but decoupled execution. This design enables fine-grained responsiveness while maximizing intra-GPU parallelism.
@@ 8. Conclusion
While the asymmetric resource demands of prefill and decode phases in LLM inference is well recognized, its implications for intra-GPU coordination have remained underexplored. In this work, we analyze how co-executing these phases under coarse-grained batching strategies leads to resource imbalance and performance interference. To address this, we propose Nexus, a novel SM-partitioned execution framework with dynamic resource allocation and phase-specific schedulers. Our system consistently improves end-to-end throughput and latency across diverse LLM workloads, outperforming state-of-the-art serving systems.